Bob works in an office of 100 people. He received a score of 4 on his evaluation. The mean score for everyone in the office is 3, the standard deviation is 1.58. What is his standardized score (z-score)?
Answer #1

Z score=(raw score-mean)/sd

Raw score=4

Mean=3

Sd=1.58

Z score=(4-3)/1.58=1/1.58=.632911